# Autonomous Horizon-based Asteroid Navigation With Observability-constrained Maneuvers

### v1

@@ 2.2 Horizon-based OpNav
The Christian-Robinson OpNav algorithm [16] and its analytical measurement covariance formulation [34] are utilized for the measurement implementation. It is an efficient and accurate computing method to calculate the position of an observer with respect to the center of mass of a target body as it directly computes from the edge-detected horizon points rather than using curve fitting. The observations for the EKF and the simulation are derived from this OpNav process.
The inputs include the pixel coordinates of all the detected points in the camera frame, the inverse camera calibration matrix, the attitude transformation matrix between the parent body and the camera, and the triaxial parameters of the shape of the target body. The inverse camera matrix is defined as follows:
@@ 2.3 Extended Kalman Filter
An Extended Kalman Filter (EKF) is used to perform the state estimation using the OpNav measurements. Testing the EKF’s performance across various operational scenarios helps establish the algorithm’s robustness and identify areas for enhancement. First, some key assumptions include the satellite’s ability to continuously observe the asteroid with its optical camera via nadir pointing (i.e., the camera is always oriented to point towards the center of the target body), the attitude of the spacecraft relative to the Hill frame is always known, the asteroid shape is known, and the availability of initial state estimates with inherent uncertainties. These assumptions are made to simplify the problem and focus on estimating the position. Constraints involve the limitations of the camera’s resolution and the accuracy of onboard sensors, the maximum semi-major axis for a stable orbit in this system, and the computational complexity of generating and processing synthetic observation images in real time. The EKF algorithm used is as follows:
@@
2.4 Lyapunov Control
Lyapunov control is a method that ensures dynamical system stability by developing Lyapunov functions in a nonlinear feedback control case. It is beneficial as it allows for creating an analytical controller with light computational cost and is easy to implement. Furthermore, it is possible to incorporate path constraints using artificial penalty functions. However, the Lyapunov controller does not guarantee satisfaction of the path constraint and requires parameter tuning. It may also not return an optimal solution, but its efficiency and satisfactory performance prove it to be a good option for this research. A Lyapunov function [35] is a scalar continuous function $V(\bm{x})$ that is locally positive definite about a reference state $\bm{x}_{r}$ of an autonomous system $\dot{\bm{x}}=\bm{f}(\bm{x})$.
In spacecraft control, Lyapunov functions can be designed to guide the spacecraft’s state toward a desired configuration, such as a stable orbit, attitude, or particular rendezvous point [36]. By defining an appropriate Lyapunov function for a spacecraft’s position and velocity, the system can be controlled to remain stable under perturbations, such as SRP or irregular gravitation in the case of an asteroid. control inputs are applied to approach a desired target state. These inputs are derived from the gradient of the Lyapunov function, ensuring that the spacecraft moves in a direction that reduces the overall system energy. The approach is to design a controller $\bm{u}=\bm{c}(\bm{x})$ such that $V(\bm{x})$ is a Lyapunov function.
3 Synthetic Image Generation Pipeline
The image generation is done using built-in Matlab graphics tools. The procedure involved illuminating a white sphere, ellipsoid, or 3-dimensional mesh model of an asteroid against a black background with the mean radius of Bennu (or any sample asteroid) using an infinitely far away light source that is aligned to originate from the negative $x$-axis in the Hill frame to represent a parallel ray light source like the Sun. The camera is then placed based on an input distance, azimuth, and elevation calculated from the current spacecraft position and an image is captured. Gaussian white noise is added to the image to add realistic complexity for the filter to process. Edge detection is executed by identifying high gradient magnitudes between pixels on the image, which is commonly where the lit limb edge meets blank space. The generated images were compared with those produced by existing higher fidelity image generators in identical conditions and found to be sufficiently accurate when comparing the resulting OpNav measurement. Below is a stepwise summary of the process utilized for the synthetic image generation:
  1. 1.
    Generate a sphere, ellipsoid, or mesh scaled according to the camera settings and distance or an experimentally determined scale factor. A fixed scale factor of 1.3065 was experimentally found to produce the best OpNav performance with minimal measurement bias to adjust according to the inherent scaling bias encountered when generating images using the MATLAB figure environment. This was found by visually comparing generated images with real images and verifying the OpNav measurement across a range of distances, then fitting a scale factor based on the linear relationship between error and distance.
  2. 2.
    Rotate the object using a specified axis and angle. In the context of this paper, Bennu’s retrograde rotation properties are used. The axis is at approximately 180 [37] from the z-axis in the Hill frame with a synodic rotation period of 4.296057 hours [38]. This information is used to calculate and apply the angle by which the object had rotated at each time step during simulation with the equation $\theta_{\textrm{asteroid}}=\frac{t}{t_{\textrm{rp}}}360^{\circ}$ where $\theta_{\textrm{asteroid}}$ is the angle by which the asteroid has rotated, $t$ is the time elapsed and $t_{\textrm{rp}}$ is the asteroid’s synodic rotation period.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
    Set the figure size and axis limits according to the image size, apply the camera’s FOV setting to the axes’ camera view angle, and set aspect ratios to be equal
  7. 7.
The synthetic image generation process is also expressed in Algorithm 1:
Algorithm 1 Asteroid Image Generation Procedure
1:procedure $img$=ImageGenerator($x\_y\_z,up,dist,FOV,imgSize,rad,ellip\_shape,R\_ellip$)
2:Input:
3:     $x\_y\_z$: Camera direction vector
4:     $up$: Camera ”up” pointing direction vector
5:     $dist$: Distance from camera to object center
6:     $FOV$: Camera field of view
7:     $imageSize$: Desired square image dimension in pixels
8:     $rad$: Characteristic radius of the asteroid
9:     $ellip\_shape$: Scaling factors $(a,b,c)$ matrix defining ellipsoid shape
10:     $R\_ellip$: Rotation angle about the chosen rotation axis
11:     $R\_axis$: Rotation axis vector in body frame
12:Output:
13:     A grayscale image of the half-illuminated ellipsoid approximating the asteroid.
14:     Generate object (sphere, ellipsoid or mesh) $[X,Y,Z]$ coordinates using center=(0,0,0)000(0,0,0)( 0 , 0 , 0 ) and $ellip\_shape$
15:     Scale object by a scale factor of 1.3065
16:     Rotate Object by $R\_axis$ and angle=$R\_ellip$
17:     Create a figure with a black background and invisible, equal, and tight axes
18:     Plot object with invisible edges and white color
19:     Set viewDirection=$x\_y\_z$
20:     Set cameraPositionNormalize=$dist\cdot x\_y\_z$
21:     Set cameraUpDirection=$up$
22:     Set light at infinite distance with direction=$[-100]$ and Gouraud type
23:     Set dull material with 0 ambient and specular, full diffuse reflectivity and shininess coefficient of 10
24:     Set figure size = $imgSize$ and axis limits = [$-imgSize/2$,$imgSize/2$]
25:     Apply camera field of view = $FOV$
26:     Convert image to grayscale
27:     Apply noise to the image
28:     Return the image
29:end procedure
A few examples of generated images are shown in Fig. 2
Figure 2: Example synthetic images of a half-illuminated sphere with noise (left), ellipsoid (center), and Bennu model at a distance of 5 radii (right)
A gradient-based approach detects the horizon points from the generated synthetic images. First, calculate the gradient of the image matrix and find the gradient magnitude at each pixel. Next, find the maximum gradient magnitude across the entire image. Then, multiply this value by 0.7 to define a minimum gradient threshold and find the pixel coordinates of all points with a gradient magnitude above it. These are the desired horizon points. The value of 0.7 was chosen for the best performance after some experimental testing. To avoid detecting random noisy pixels, the absolute minimum gradient threshold was set to 0.1, which was determined experimentally by finding an appropriate threshold above the maximum gradient in dark-side lighting conditions but below the maximum gradient of regular lighting conditions. The edge detection process is expressed in Algorithm 2:
Algorithm 2 Gradient Edge Detection Procedure
procedure $points$=GradientEdgeDetect($img,threshold$)
2:Input:
     $img$: A grayscale image of the half-illuminated ellipsoid approximating the asteroid.
4:     $threshold$: The gradient threshold used to determine edge points.
Output:
6:     $points$: A $3\times n$ matrix containing the coordinates of the lit edge pixels detected with each pixel stored in the form $[x,y,1]^{\top}$
     Convert $img$ to numerical double format ($img\_dbl$) for gradient calculation
8:     Calculate the 2D $gradient$ of $img\_dbl$
     Compute the $gradient\_magnitude$ from the $gradient$
     Add all points with $gradient\_magnitude>minimum\_gradient$ to $points$
12:     Return $points$
end procedure
@@ 4 Observability-constrained Controller Design
The goal is to design a controller that generates a controlled trajectory along which the view of the asteroid is always in favorable lighting conditions. The objective is to develop a Lyapunov controller that avoids certain regions where the OpNav algorithm performs poorly. These poor observability regions have been characterized experimentally in Section 5.
@@ 4.1 Path Constraint Formulation
    Minimum radius constraint: $R_{\min}-\left\lVert{}^{H}\bm{r}\right\rVert_{2}\leq 0$ with $R_{\min}=2R_{\mathrm{asteroid}}$
@@
    Maximum radius constraint: $\left\lVert{}^{H}\bm{r}\right\rVert_{2}-R_{\max}\leq 0$ with $R_{\max}=25R_{\mathrm{asteroid}}$
@@
where $R_{\textrm{asteroid}}$ is the radius of the asteroid.
@@
$h_{1}$ is the component of angular momentum in the x-axis direction in the Hill frame. The partial derivatives of the path constraints are computed in appendix A
@@ 4.2 Controller Derivation
Milankovitch elements are used for the controller because they have no singularity and can target a certain orbit state over a specific point. We start by only targeting the slow variables and freeing the fast variable, which in this case would be the true longitude. Although the rotation between the Hill and inertial frames was found to affect the angular momentum and eccentricity vector components of $\bm{f}_{0}$ in Section 2.1, it is possible to assume $\bm{f}_{0,slow}\approx 0$. Therefore, ${}^{H}\dot{\bm{x}}_{\textrm{slow}}=B_{\textrm{slow}}{{}^{H}}\bm{u}$. This assumption is reasonable because the orbital angular velocity of Bennu is approximately $(360/436.649)=0.824\mathrm{[degrees/day]}$ or $1.665\times 10^{-7}\mathrm{[rad/s]}$ which only leads to a total rotation of about 5 degrees over the duration of the simulations in this research, which is at most 6 days long. For reference, 436.649 days is the orbital period of Bennu [38]. However, this assumption is only applied for the controller derivation and only affects the theoretical stability guarantees in the context of Lyapunov control. Numerical dynamics propagations are based on the true dynamics. Thus, the error state can be defined as follows:
@@
where $w$ is the weight of the penalty, and $P$ is the penalty function in terms of the path constraint $g(^{H}\bm{x})\leq 0$, which must be negative at the target state and smooth everywhere. The penalty function should also monotonically increase in $g$ for $g>0$. For this, we define the penalty functions as set to 00 when the constraint condition is not violated and greater than 00 when it is. This can be represented as follows:
@@
5 OpNav and EKF Performance Analysis
To start, a simplified base case is used by choosing the asteroid Bennu for which abundant information is available [2, 39]. This is ideal for testing and tuning the filter as simulated ’truth’ and filter predictions can be validated. The relevant initial conditions and properties used are shown in Table 1, where $M$ is the asteroid’s mass, $e$ is the eccentricity of the asteroid’s orbit, and $r$ is the mean radius of the asteroid.
@@ 5 OpNav and EKF Performance Analysis
Parameter Value
$M$ $7.329\times 10^{11}\,\text{kg}$
$\mu$ $4.890\times 10^{-9}\,\text{km}^{3}/\text{s}^{2}$
$\mu_{\text{Sun}}$ $1.327\times 10^{1}1\,\text{km}^{3}/\text{s}^{2}$
$G_{1}$ $1\times 10^{8}\,\text{kg.km}^{3}/(\text{s}^{2}.\text{m}^{2})$
$B$ $33$
$a$ $1.685\times 10^{8}\,\text{km}$
$e$ $0.204$
$d$ $1.720\times 10^{8}\,\text{km}$
@@
Additionally, the OpNav parameters used for this research are shown in Table 2.
@@
Parameter Value
$\theta_{\text{FOV}}$ $90^{\circ}$
$S$ 1000 pixels
$\alpha$ 0
$dx$ 1
$dy$ 1
5.1 Distance and Angle Accuracy Analysis
The OpNav algorithm’s accuracy at varying distances and lighting conditions is examined by running multiple simulations with noise at varied fixed points.
@@ 5.1 Distance and Angle Accuracy Analysis
As can be seen in Fig. 3, the OpNav algorithm is highly effective at low range but starts to involve major error at distances greater than 25-30 times the radius of the asteroid. In this case, we can see that deviation becomes visibly significant after a distance of around 10 km. However, the analytical covariance still provides a good envelope for the errors. We can define one range observability constraint based on this result. While OpNav still provides results beyond this, eventually, the magnitude of the error exceeds the absolute value of the distance itself, making it unusable after a certain point. Therefore, the algorithm must be used within this range, which has now been successfully determined. This distance is still a greater capability than near-range observation methods, which require the spacecraft to be much closer than 25-30 times the asteroid’s radius.
@@
We can infer from Fig. 4 that the algorithm is accurate for most angles except straight-on due to the terminator effect and angles close to the dark side as the size of the lit horizon shrinks. The performance deteriorates if the spacecraft views the asteroid from within 30 degrees of the $+x$ axis near the dark side. The covariance is undefined when viewing the dark side, as no horizon points are detected. The error is accordingly equivalent to the true distance of the spacecraft in the implementation used for this research, as the OpNav algorithm returns a zero measurement due to an absence of input horizon points. In other words, the measurement is unusable in these lighting conditions. From this experiment, we define a keep-out cone to ensure observability, which is given as a cone with a 30-degree half-angle along the positive $x$-axis.
5.2 Elliptic Shape Performance
We now test the OpNav performance against increasingly elliptical bodies. The shape ratios are represented using the notation [a,b,c] where a, b and c represent the principal axis dimensions of the triaxial ellipsoid body and the principal axis frame is defined such that $a\geq b\geq c$ according to convention [16]. To isolate the effect of body ellipticity, rotation is ignored for this test. As can be seen in Fig. 5, the OpNav error and covariance increase with highly elliptic bodies. The error is correctly bounded by the covariance for all cases up to an ellipsoid with a shape ratio of [5,1,1], supporting its usability for state estimation with ellipsoid bodies. For this test case, we simulate along one revolution of an FTO with a radius of $2.0429$km.
5.3 Rotating Ellipsoid Performance
The elliptic shape scenario is repeated with rotation to determine its effect on OpNav performance. A [2.5,1,1] ellipsoid with the rotation period and axis of Bennu as described in Section 3 is used. For this test case, we simulate along one revolution of an FTO with a radius of $2.0429$km. The result suggests that rotation significantly impacts the accuracy of OpNav, likely due to various alignments leading to lit limb geometries that are better or worse for accuracy. While the measurement error was found to be correctly bounded in the non-rotating [5,1,1] ellipsoid test case, the [2.5,1,1] ellipsoid seems to be near the limit of usability when considering rotating bodies. The error is marginally bounded by the covariance in this case and the error magnitude is significant.
5.4 Edge Detection Threshold Analysis
This implementation uses a pixel gradient method to detect the asteroid’s lit limb edge. The maximum gradient on the image is found. All pixels with a gradient within a certain threshold of the maximum gradient are stored as the detected edge points for the OpNav measurement. The selection of this gradient, a value between 0 and 1, is critical to the measurement performance. Thus, we analyzed the performance for varying thresholds. The difference between the limb detected by a high threshold and the ideal threshold is shown in Fig. 7. Lower thresholds were found to lead to a higher lit limb angle and number of detected points, coupled with a lower analytical covariance. Using a lower threhshold may also increase the risk of misdetecting background noise or excess points (such as those along the terminator line) leading to inconsistencies between the analytical and measurement covariance. The opposite effect holds for higher thresholds. Its error did not exceed the measurement covariance as lower thresholds may have due to detecting an excess number of points or including background noise. The ideal balance was found to be a value of 0.4 as it had consistently improved performance from higher thresholds, particularly for ellipsoid cases. At this threshold, sufficient points are included to characterize the asteroid silhouette, while avoiding inconsistencies in analytical and measured covariances or detecting erroneous points.
[…]
5.5 Trajectory Testing
The OpNav algorithm and EKF are tested across various types of trajectories, such as hyperbolic approaches and flybys, distant encirclements, and more, to ascertain properties of its performance across various situations and deduce when it is more effective or may have room for further improvement.
These initial conditions are arbitrarily chosen in terms of both position and velocity to provide some variation and generate natural paths that a spacecraft may follow depending upon its nature of approach to an asteroid. A few sample results are in Fig. 8. The first case depicted a hyperbolic flyby from a large distance. The second test is conducted at a medium range, keeping a similar distance from the asteroid while changing the lighting condition. The last case demonstrates another flyby where the asteroid is approached closely, and the spacecraft passes through the dark side.
Figure 8: Sample OpNav Testing Trajectories with a hyperbolic approach (top left), distant flyby (top right) and single revolution close proximity flyby (bottom)
Based on the behavior in Fig. 8, we can infer that position estimation is best when approaching near-linearly from the same angle or in close proximity to the asteroid. Horizon-based OpNav, and hence EKF, become inaccurate when the angle varies extremely in a short time or has a highly nonlinear trajectory as seen in the divergence of the OpNav measurements in each test case. Horizon-based OpNav performs well within medium range when in a hyperbolic approach trajectory before the view angle starts changing rapidly (before/after the viewpoint).

### v2

@@ 2.2 Horizon-based OpNav
The CRA [16] and its analytical measurement covariance formulation [56] are utilized for the measurement implementation. The observations for the EKF and the simulation are derived from this OpNav process. The measurements are weighed by the analytical measurement covariance which propagates through the standard deviation of the detected horizon points $\sigma_{\mathrm{pix}}$ and is directly used as the measurement noise covariance $R_{k}$ in the EKF’s update step. Since we use synthetic image generation for this research, $\sigma_{\mathrm{pix}}$ is approximated by representing the detected limb points in polar form relative to the image center then computing the standard deviation of the radii scaled for ellipticity as outlined in previous studies that apply OpNav to synthetic images [57, 58]. The inputs include the pixel coordinates of all the detected points in the camera frame, the inverse camera calibration matrix, the attitude transformation matrix between the parent body and the camera, and the triaxial parameters of the shape of the target body. The inverse camera matrix is defined as follows:
@@
The known attitude assumption has two types of consequences. First, in terms of relative small body-camera attitude determination, the target small body’s inertial attitude can be determined accurately by determining its pole direction and rotation rate through a variety of methods [59, 60] such as the recently developed PoleStack method that can determine the target’s rotational motion onboard the spacecraft as it approaches the small body which allows definition of a rotating, target-fixed reference frame and in turn the small body inertial attitude [61]. The spacecraft’s inertial attitude is known from star trackers [62, 63] and the camera’s orientation on the spacecraft is known. These combined usually give an attitude error less than a degree [16, 64]. On the other hand, attitude control error is effectively thruster directional error, which are also generally on the scale of a degree or lower based on previous small body missions [65, 66]. Therefore, known attitude is a reasonable assumption in the context of this research.
@@ 2.3 Extended Kalman Filter
An Extended Kalman Filter (EKF) is used to perform the state estimation using the OpNav measurements. Testing the EKF’s performance across various operational scenarios helps establish the algorithm’s robustness and identify areas for enhancement. First, some key assumptions include the satellite’s ability to continuously observe the small body with its optical camera via nadir pointing (i.e., the camera is always oriented to point towards the center of the target body), the attitude of the spacecraft relative to the Hill frame is always known, the small body shape is known, and the availability of initial state estimates with inherent uncertainties. These assumptions are made to simplify the problem and focus on estimating the position. Constraints involve the limitations of the camera’s resolution and the accuracy of onboard sensors, the maximum semi-major axis for a stable orbit in this system, and the computational complexity of generating and processing synthetic observation images in real time. The EKF algorithm used is as follows:
@@
[…]
We note that for all figures depicting distance norm error from either the OpNav measurement or EKF estimate in this paper, we use the square root of the maximum eigenvalue of the position covariance ($3\sqrt{\lambda_{max}(P_{\bm{r}})}$) to define the error bounds.
@@ 3 Observability-constrained Controller Design
In spacecraft control, Lyapunov functions can be designed to guide the spacecraft’s state toward a desired configuration, such as a stable orbit, attitude, or particular rendezvous point [67]. By defining an appropriate Lyapunov function for a spacecraft’s position and velocity, the system can be controlled to remain stable under perturbations, such as SRP or irregular gravitation in the case of an small body. Control inputs are applied to approach a desired target state. These inputs are derived from the gradient of the Lyapunov function, ensuring that the spacecraft moves in a direction that reduces the overall system energy. The approach is to design a controller $\bm{u}=\bm{c}(\bm{x})$ such that $V(\bm{x})$ is a Lyapunov function.
While Lyapunov-based controllers have been extensively studied for spacecraft orbit and attitude stabilization around small bodies [50, 51, 49], prior works generally focus on stability with respect to gravitational and dynamical perturbations, without explicitly incorporating real-time navigation performance metrics into the control design. This section develops an observability-constrained Lyapunov controller in which the penalty functions are derived directly from experimentally measured poor performance regions of a horizon-based OpNav algorithm (CRA). These poor observability regions have been characterized experimentally in Section 4. Unlike conventional constraint handling in Lyapunov control which often uses generic keep-out zones or safety envelopes, our formulation parameterizes the path constraints in closed form using OpNav performance and observability limitations then augments the Lyapunov function with smooth, continuously differentiable artificial potentials that are compatible with stability proofs while being computationally lightweight. This combined integration of navigation observability consideration within small body dynamical environments with artificial potential constraints results in a feedback law that both maintains stability and actively steers trajectories away from poor-visibility regions while being real-time onboard feasible. By targeting slow variables and isolating them from fast orbital phase effects, the derivation also reduces onboard computational load without sacrificing robustness. Evidence of constraint satisfaction by the Lyapunov controller derived in this paper is included with the numerical results in Section˜5.
@@ 3.1 Path Constraint Formulation
    Minimum radius constraint: $R_{\min}-\left\lVert{}^{H}\bm{r}\right\rVert_{2}\leq 0$ with $R_{\min}=2R_{\mathrm{sb}}$
@@
    Maximum radius constraint: $\left\lVert{}^{H}\bm{r}\right\rVert_{2}-R_{\max}\leq 0$ with $R_{\max}=25R_{\mathrm{sb}}$
@@
where $R_{\textrm{sb}}$ is the radius of the small body.
@@
$h_{1}$ is the component of angular momentum in the x-axis direction in the Hill frame. The partial derivatives of the path constraints are computed in appendix˜C
@@ 3.2 Controller Derivation
We start the controller derivation by only targeting the slow variables and freeing the fast variable, which in this case would be the true longitude. Although the rotation between the Hill and inertial frames was found to affect the angular momentum and eccentricity vector components of $\bm{f}_{0}$ in Section˜2.1, it is possible to assume $\bm{f}_{0,slow}\approx 0$. Therefore, $^{H}\dot{\bm{x}}_{\textrm{slow}}=B_{\textrm{slow}}{{}^{H}}\bm{u}$. This assumption is reasonable because the orbital angular velocity of the small body (assuming Bennu’s properties) is approximately $(360/436.649)=0.824\mathrm{[degrees/day]}$ or $1.665\times 10^{-7}\mathrm{[rad/s]}$ which only leads to a total rotation of about 5 degrees over the duration of the simulations in this research, which is at most 6 days long. For reference, 436.649 days is the orbital period of the small body considered [68]. However, this assumption is only applied for the controller derivation and only affects the theoretical stability guarantees in the context of Lyapunov control. Numerical dynamics propagations are based on the true dynamics. Thus, the error state can be defined as follows:
@@
where $w$ is the weight of the penalty, and $P$ is the penalty function in terms of the path constraint $g(^{H}\bm{x})\leq 0$, which must be negative at the target state and smooth everywhere. The penalty function should also monotonically increase in $g$ for $g>0$. For this, we define the penalty functions as set to 0 when the constraint condition is not violated and greater than 0 when it is. This can be represented as follows:
@@
4 Determining Poor Observability Regions
This section demonstrates the method of determining observability-constraining path constraints for the controller, so that the derivation and approach can be generalized and applied to other types of optical navigation methods. All simulations in this paper are conducted using only spheres and triaxial ellipsoids using the radius and shape ratio parameters of the asteroid Bennu for which abundant information is available [2, 55]. We note that this is a simplified base case ideal for testing and tuning the filter and controller as simulated ’truth’ and filter predictions can be validated with realistic small body physical properties and dynamical scenario, which aligns with the primary control focus of this research. The CRA cannot be directly applied to an irregular body such as Bennu without further extension, which is a problem not considered or solved in this paper and instead suggested as a step for future research [69]. reiterate that this paper does not extend image generation or processing, or the CRA for applications to Bennu or other irregular bodies. This is The relevant initial conditions and properties used are shown in Table˜1, where $M$ is the small body’s mass, $e$ is the eccentricity of the small body’s orbit, and $r$ is the mean radius of the small body.
@@ 4 Determining Poor Observability Regions
Parameter Value Unit Description
$M$ $7.329\times 10^{11}$ kg Small body mass
$\mu$ $4.890\times 10^{-9}$ km3/s2 Small body gravitational parameter
$\mu_{\text{Sun}}$ $1.327\times 10^{11}$ km3/s2 Sun gravitational parameter
@@
$m/A$ $33$ kg/m2 Spacecraft mass-to-area ratio
$a$ $1.685\times 10^{8}$ km Small body orbital semi-major axis
$e$ $0.204$ Small body orbital eccentricity
$d$ $1.720\times 10^{8}$ km Heliocentric distance at epoch
$r$ $0.241$ km Small body mean radius
Additionally, the OpNav parameters used for this research are shown in Table˜2.
@@
Parameter $\theta_{\text{FOV}}$ $S$ $\alpha$ dxdx dydy
Value $30^{\circ}$ 1000 pixels 0 1 1
These settings are chosen based on camera settings used on real space missions [3, 4, 5, 6, 14, 65, 66, 64, 70, 71]. An intermediary FOV value is chosen to allow the largest range of capabilities. A narrow FOV camera allows resolution at farther distances but leads to the object exceeding the image frame in proximity, while a wide FOV camera only leads to resolved images at closer distances which may be too small from a safety perspective. Thus, a range of 10-30 is generally ideal for most applications. The image size matches realistic cameras and the other parameters are chosen to emulate an ideal camera since modeling and testing non-ideal camera parameters is outside the focus of this research.
The OpNav algorithm’s accuracy at varying distances and lighting conditions is examined by running multiple simulations with noise at varied fixed points. A perfectly spherical target body is used for this analysis as we intend to determine an exact observability cutoff in the ideal case.
4.1 Distance Accuracy Analysis
@@
As previously mentioned, for all figures depicting distance norm error from either the OpNav measurement or EKF estimate in this paper, we use the square root of the maximum eigenvalue of the position covariance ($3\sqrt{\lambda_{max}(P_{\bm{r}})}$) to define the error bounds.
@@
As can be seen in Fig.˜2, the OpNav algorithm is highly effective at low range but starts to involve major error at distances greater than 25 km, in this case more than 100 times the radius of the small body for the combination of $\theta_{FOV}$ and $S$ in Table˜2. This is because as range increases, the small body’s apparent radius in pixels shrinks and the measurement becomes progressively less informative. However, the analytical covariance still provides a good envelope for the errors. We can define one range observability constraint based on this result. We may define a practical usable "observability limit" when the error bound reaches $10\%$ of the range, which is close to the aforementioned 25 km threshold as seen in Fig.˜2. While OpNav still provides results beyond this, eventually, the magnitude of the error exceeds the absolute value of the distance itself, making it unusable after a certain point. Therefore, the algorithm must be used within this range, which has now been successfully determined.
Importantly, this threshold is not a fixed distance. it depends on camera intrinsics and the target’s physical radius. For a given camera, a larger body remains observable to greater distances, while a smaller body reaches the same ill-conditioning sooner. Moreover, when the object’s image size varies very slowly with distance (far-field), multiple nearby ranges can yield effectively indistinguishable silhouettes, leading to similar solutions over a band of distances. This explains the flattening/ambiguity observed at long range and motivates a distance (or apparent-size) cutoff for reliable autonomous operation.
4.2 Angle Accuracy Analysis
@@
We can infer from Fig.˜3 that the algorithm is accurate for most angles except straight-on due to the terminator effect and angles close to the dark side as the size of the lit horizon shrinks. The performance deteriorates if the spacecraft views the small body from within 30 degrees of the +x+x axis near the dark side. The covariance is undefined when viewing the dark side, as no horizon points are detected. The error is accordingly equivalent to the true distance of the spacecraft in the implementation used for this research, as the OpNav algorithm returns a zero measurement due to an absence of input horizon points. In other words, the measurement is unusable in these lighting conditions. From this experiment, we define a keep-out cone to ensure observability, which is given as a cone with a 30-degree half-angle along the positive $x$-axis in the Hill frame.